OpenStudy (anonymous):

Solve the exponential equation 7(2^5x-6) = 56

OpenStudy (lgbasallote):

divide both sides by 7 first...what do you get?

OpenStudy (anonymous):

2^5x-6 = 8

OpenStudy (anonymous):

I think I see it now. then it's 2^5x-6 = 2^3

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

now equate the exponents

OpenStudy (anonymous):

ahh got it, thanks!

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

is the answer 9/5? that's what I got

OpenStudy (lgbasallote):

yup

OpenStudy (unklerhaukus):

you got it